OpenStudy (anonymous):

I guys i have a question on this COMPLETING THE SQUARE : solve by completing the square; y=6x^2-7x+2

OpenStudy (anonymous):

yup your gonna have to complete the square and work with some fractions

OpenStudy (anonymous):

i plugged it in the calculator the answer is 38 y=38

OpenStudy (anonymous):

so far i am stuck because

OpenStudy (agreene):

\[(x-\frac{7}{12})^2=\frac1{144}\] is your completed square. \[x=\frac12; \space x=\frac23\]

OpenStudy (anonymous):

okay this is what i have so far: 0=6x2-7x+2 -2=(7/2)^2=6x^2-7x+9(7/2)^2

OpenStudy (anonymous):

do you have a scientific calculator ?

OpenStudy (anonymous):

no i have to show my work

OpenStudy (anonymous):

and I need help on HOW you get there, not just the answer

OpenStudy (agreene):

I rather hate typing out steps with fractions--but wolfram did it just as I did. Click show steps on the right. http://www.wolframalpha.com/input/?i=0%3D6x%5E2-7x%2B2+solve

OpenStudy (anonymous):

what is the perfect share of 6x^2-7x+49/4

OpenStudy (anonymous):

* sqare

OpenStudy (anonymous):

(49/4)

OpenStudy (agreene):

that isnt going to factor to something useful.

OpenStudy (anonymous):

i know that's why i am stuck

OpenStudy (anonymous):

no i know how to do it but that just stumped me

OpenStudy (agreene):

Well, if you click the link I gave, and then show steps under results... you'll see the way I did it. I just seriously hate typing fractions on here.

OpenStudy (anonymous):

okay dude i knpow how to do completing the squares in general

OpenStudy (agreene):

and 6x^2-7x+49/4 factors to: 1/4 (24 x^2-28 x+49)

OpenStudy (anonymous):

JUST ! stumped me because there is no perfect sqare for 6x^@-7x+(49/4)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are going to have to divide by6 as step one. you cannot complete the square if the leading coefficient is not 1

OpenStudy (anonymous):

humm?

OpenStudy (anonymous):

also i note that this says solve by completing the square; y=6x^2-7x+2 this is not clear. you can "solve" \[6x^2-7x+2=0\] or you can "complete the square" and rewrite \[y=6x^2-7x+2\] as \[y=6(x-\frac{7}{12})^2-\frac{1}{24}\]

OpenStudy (anonymous):

where did the 12 and 24 xcome from

OpenStudy (anonymous):

if the question is "solve" \[6x^2-7x+2=0\] by completing the the square you are justified in dividing by 6 because you have an equation to solve. start with \[6x^2-7x=-2\] then \[x^2-\frac{7}{6}x=-\frac{1}{3}\]

OpenStudy (anonymous):

i thought it would be 2/3

OpenStudy (anonymous):

on the right side of the 7/6 thing

OpenStudy (anonymous):

now half of \[\frac{7}{6}\] is \[\frac{7}{12}\] so next step is \[(x-\frac{7}{12})^2=-\frac{1}{3}+\frac{49}{144}\]

OpenStudy (anonymous):

now a step of arithmetic gives \[(x-\frac{7}{12})^2=\frac{1}{144}\] and then \[x-\frac{7}{12}=\frac{1}{12}\] or \[x-\frac{7}{12}=-\frac{1}{12}\]

OpenStudy (anonymous):

finally \[x=\frac{8}{12}=\frac{2}{3}\] or \[x=\frac{6}{12}=\frac{1}{2}\]

OpenStudy (anonymous):

and that is all the gory complete square details

OpenStudy (anonymous):

okay ummm thanks

OpenStudy (anonymous):

if not clear let me know but i think all the details are written out without any thing missing

OpenStudy (anonymous):

yw